OpenStudy (anonymous):

We have a small deck of ten cards. Five of the cards are red, three are blue and two are green. We randomly select four cards from the deck with replacement. That is, after we select a card and record the colour, we put the card back in the deck and thoroughly shuffle them before we select another card.

OpenStudy (anonymous):

make a tree diagram

OpenStudy (anonymous):

how should i make a tree diagram for this probability?

OpenStudy (nincompoop):

http://bit.ly/141H2wI

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

What is the probability that the rst two selected cards are the same colour?

OpenStudy (anonymous):

first

OpenStudy (anonymous):

are ut here?

OpenStudy (anonymous):

hello?

OpenStudy (kropot72):

Did the full question ask you to make a tree diagram?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What is the probability that the rst two selected cards are the same colour?

OpenStudy (anonymous):

thats the question

OpenStudy (anonymous):

What is the probability that the first two selected cards are the same colour?

OpenStudy (kropot72):

There are three possible events to consider: 1. P(red + red) = 5/10 * 5/10 2. P(blue + blue) = 3/10 * 3/10 3. P(green + green) = 2/10 * 2/10 These three events are mutually exclusive, therefore the probability of the first two selected cards being the same color is the sum of the above three probabilities.

OpenStudy (kropot72):

@yashar806 Have you done the calculation yet?

OpenStudy (kropot72):

Required probability is \[0.5^{2}+0.3^{2}+0.2^{2}=you\ ca\ calculate\]

OpenStudy (kropot72):

you can*

OpenStudy (anonymous):

yep